# Selective Attention Federated Learning: Improving Privacy and Efficiency for Clinical Text Classification

### v1

@@ V-C Layer Selection Analysis
We analyze the distribution of selected layers across different tasks and clients. Figure LABEL:fig:layers shows that SAFL consistently identifies and updates task-critical layers, with higher layers (closer to the output) being selected more frequently for concept extraction, while middle layers are more important for diagnosis code prediction.
@@ VI Conclusion
Our work pushes the state-of-the-art in federated clinical NLP by demonstrating that an attention-informed FL paradigm can achieve strong privacy preservation and efficiency concurrently. As federated training of language models in healthcare continues to attract attention, we provide a timely step towards frameworks that do not force a compromise between privacy and model quality. The selective dynamics in our approach share theoretical foundations with control systems work, such as point regulation [27] and boundary control methods [28], where targeted resource allocation leads to more robust overall system performance. We envision future research building on this foundation in several ways. First, formal privacy guarantees (e.g., differential privacy or secure multi-party computation) could be combined with selective attention FL to further strengthen protection of patient data. Second, evaluating our method on large-scale federated biomedical language models will be important to confirm its scalability and robustness. Finally, beyond text classification, selective attention mechanisms could benefit other healthcare tasks in FL such as named entity recognition or summarization of clinical notes, where identifying critical information while ignoring sensitive details is paramount. We believe that Selective Attention Federated Learning opens up a promising direction for safe and effective collaborative learning in the medical domain.

### v3

@@ Abstract
Index Terms:
Federated Learning, Natural Language Processing, Healthcare, Privacy, Selective Attention
**footnotetext: Equal contribution
@@ V-C Layer Selection Analysis
We analyze the distribution of selected layers across different tasks and clients. SAFL consistently identifies and updates task-critical layers, with higher layers (closer to the output) being selected more frequently for concept extraction, while middle layers are more important for diagnosis code prediction.
@@ VI Conclusion
Our work pushes the state-of-the-art in federated clinical NLP by demonstrating that an attention-informed FL paradigm can achieve strong privacy preservation and efficiency concurrently. As federated training of language models in healthcare continues to attract attention, we provide a timely step towards frameworks that do not force a compromise between privacy and model quality. The selective dynamics in our approach share theoretical foundations with control systems work, such as point regulation [cai2025set] and boundary control methods [cai2025inverse], where targeted resource allocation leads to more robust overall system performance. We envision future research building on this foundation in several ways. First, formal privacy guarantees (e.g., differential privacy or secure multi-party computation) could be combined with selective attention FL to further strengthen protection of patient data. Second, evaluating our method on large-scale federated biomedical language models will be important to confirm its scalability and robustness. Finally, beyond text classification, selective attention mechanisms could benefit other healthcare tasks in FL such as named entity recognition or summarization of clinical notes, where identifying critical information while ignoring sensitive details is paramount. We believe that Selective Attention Federated Learning opens up a promising direction for safe and effective collaborative learning in the medical domain.